OpenStudy (anonymous):

will give medals factor 2g^3 - g^2 - 8g + 4

OpenStudy (anonymous):

(g+2)(g-2)(2g-1)

OpenStudy (anonymous):

\[g^2\left( 2g-1 \right)-4\left( 2g-1 \right)=\left( 2g-1 \right)\left( g^2-4 \right)\] =(2g-1)(g+2)(g-2)

OpenStudy (anonymous):

thats not one of my option

OpenStudy (anonymous):

2nd option